# Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning

### v1

@@ Abstract
The emergence of large language models (LLMs) has opened up unprecedented possibilities for automating complex tasks that are often comparable to human performance. Despite their capabilities, LLMs still encounter difficulties in completing tasks that require high levels of accuracy and complexity due to their inherent limitations in handling multifaceted problems single-handedly. This paper introduces "Smurfs", a cutting-edge multi-agent framework designed to revolutionize the application of LLMs. By transforming a conventional LLM into a synergistic multi-agent ensemble, Smurfs enhances task decomposition and execution without necessitating extra training. This is achieved through innovative prompting strategies that allocate distinct roles within the model, thereby facilitating collaboration among specialized agents. The framework gives access to external tools to efficiently solve complex tasks. Our empirical investigation, featuring the mistral-7b-instruct model as a case study, showcases Smurfs’ superior capability in intricate tool utilization scenarios. Notably, Smurfs outmatches the ChatGPT-ReACT in the ToolBench I2 and I3 benchmark with a remarkable 84.4% win rate, surpassing the highest recorded performance of a GPT-4 model at 73.5%. Furthermore, through comprehensive ablation studies, we dissect the contribution of the core components of the multi-agent framework to its overall efficacy. This not only verifies the effectiveness of the framework, but also sets a route for future exploration of multi-agent LLM systems. Our code will be available at https://github.com/FreedomIntelligence/Smurfs/ in the future.
@@ 1 Introduction
Tool manipulation has traditionally been seen as a distinctive human characteristic, dating back approximately 2.5 million years Oakley and Museum (1972); Ambrose (2001). For large language models (LLMs), access to external tools can equip them with broader capabilities beyond their fixed language modeling knowledge. For example, the search engine API empowers ChatGPT to access real-time information Zhao et al. (2023). However, LLMs still face many challenges when attempting to use tools to solve tasks. These challenges include computational expense and a lack of adaptability to new tools Hao et al. (2024); Guu et al. (2020); Qin et al. (2024).
@@
This paper addresses the critical research problem of enhancing the problem-solving capabilities of LLMs through the adoption of a multi-agent system (MAS) framework Dorri et al. (2018); Van der Hoek and Wooldridge (2008). We posit that a MAS approach can significantly augment the efficacy of LLMs in handling tasks that require a high degree of precision, adaptability, and comprehensive knowledge integration.
ToolBench-all ToolBench-long Increase
ChatGPT-DFSDT 71.5 66.0 $-7.69\%$
GPT4-ReACT 72.0 65.1 $-9.58\%$
GPT4-DFSDT 77.5 69.8 $-9.94\%$
Mistral-Smurfs 79.5 82.1 $+3.27\%$
Table 1: Pass rate of models on the ToolBench benchmark I2 category subset with long settings and all settings.
@@
To this end, we introduce "Smurfs" an innovative MAS framework inspired by the collaborative and versatile nature of its namesake cartoon characters. The Smurfs framework is based on the principle that synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs. Each agent within the Smurfs framework is designed to perform specific sub-tasks, facilitating a more nuanced and effective approach to complex problem-solving. Our research delves into the architectural design, coordination mechanisms, and the operational dynamics of integrating specialized agents into a cohesive system. Through rigorous experimental evaluation, the Smurfs framework, utilizing the mistral-7b-instruct model Jiang et al. (2023), achieved a remarkable 84.4% win rate against the benchmark set by ChatGPT-ReACT Yao et al. (2022) on the ToolBench I2 and I3 benchmark Qin et al. (2024). This outcome not only sets a new state-of-the-art in the field, but also provides concrete evidence of the effectiveness of the multi-agent approach in enhancing LLM capabilities.
The structure of this paper is as follows: Section 2 presents the motivation for utilizing a multi-agent system. The methodology employed within the framework is detailed in Section 3. Subsequently, Section 4 provides an in-depth evaluation of the experiments conducted on Smurfs. Section 5 then reviews the techniques currently related to our work. Lastly, we summarize and conclude our findings in Section 6.
@@ 2 Motivation
2.1 Limited Context Length in a Single Model
@@
LLMs face considerable challenges when tasked with managing extensive contexts. As highlighted by Liu et al. (2024), these limitations become particularly noticeable in tasks requiring assimilation and processing of large inputs, like verbose tool documents and API responses. The situation worsens when LLMs are supplemented with external information, such as document retrieval or online searching Petroni et al. (2020); Ram et al. (2023); Mallen et al. (2022). Although numerous language models capable of handling larger contexts are emerging Dai et al. (2019); Dao et al. (2022), they often face significant performance degradation when the important information is located at some positions Liu et al. (2024); Shi et al. (2023). Moreover, within the MAS framework, the impact of extended contexts on performance remains unclear.
To measure the impact of extended contexts on the performance of LLMs in tool utilization tasks, we conducted a pilot study on the ToolBench benchmark Qin et al. (2024). Additional details can be found in Appendix A.1. We selected the samples with more than 3 steps performed by the ChatGPT-DFSDT method as a subset, called "ToolBench-long", to compare the performance difference with the full set. As demonstrated in Table 1, there is a significant decrease in the pass rate of existing frameworks when faced with tasks involving lengthy questions. This result supports the hypothesis that not only do extended contexts strain the models’ computational efficiency, but they also hinder their ability to accurately interpret and respond to the given instructions. The pilot study highlighted a major issue with current tool utilization frameworks: the excessive context length adversely impacts the models’ planning and execution capabilities when using tools.
Advantages of MAS: The pilot study highlights the need for a new approach beyond the traditional single-agent LLM model. MAS offers a promising solution by distributing tasks among specialized agents. This approach enhances memory efficiency, minimizes distractions, and allows for modular design and optimization of agents. Key benefits include: (1) Memory Efficiency: MAS manages memory better by assigning distinct segments of tasks to different agents, avoiding overload from processing lengthy contexts. (2) Reduced Distractions: Specialized agents focus on specific tasks, reducing interference from irrelevant information and improving overall performance. (3) Modular Design: The modular framework enables individual agent optimization, scalability, and adaptability to diverse tasks.
2.2 Tools Using and Planning in LLMs
Figure 2: A pilot study explores the relationship between dialogue turns and the accuracy of tool selection for LLMs. It demonstrates that long-context instructions undermine the ability of LLMs to select the right tools.
Using external tools in LLMs is a well-established method to enhance model capabilities. However, as the number of available tools grows, managing multiple tools effectively becomes a complex planning challenge Qin et al. (2024). This complexity often results in longer texts for tasks involving multiple tools. In the second pilot study, see details in Appendix A.2, we investigate the relationship between conversation turns and the model’s tool selection accuracy. Figure 2 shows that as the number of dialogue rounds increases, the model’s tool selection accuracy decreases linearly, underscoring the performance impact of longer texts in tool-involved tasks.
While numerous general MAS systems perform well, as shown in studies like Du et al. (2023); Liang et al. (2023), there’s a noticeable absence of multi-agent frameworks specifically designed for tool calling tasks. As highlighted in our two pilot studies, tool usage and planning tasks are particularly sensitive to context length and efficiency. Therefore, there’s a pressing need for an optimized multi-agent framework tailored to tool calling scenarios.
@@ 3.1 Framework Overview
Figure 1 illustrates the entire workflow for the Smurfs framework. Initially, the Planning Agent identifies the user’s complex request and breaks it down into manageable sub-tasks. Executor Agents are then tasked with collecting this specific information, utilizing access to external tools. Answer Agent compiles the findings into a cohesive response, which is subsequently verified by the Verifier Agent to ensure accuracy and relevance. This process exemplifies the framework’s capability to efficiently handle complex queries by leveraging the specialized roles of multiple agents, thereby ensuring both the precision of task execution and the quality of the output. In the following sections, the functions of each agent will be detailed. More details about the memory of each agent can be seen at Appendix B
@@
The primary responsibility of the Planning Agent is task decomposition. The strategy known as least-to-most prompting Zhou et al. (2023) is highly effective in dissecting intricate problems into manageable sub-tasks and resolving them sequentially. This approach has demonstrated significant effectiveness and broad applicability. In scenarios involving complex reasoning tasks, we utilize the fundamental principles of the least-to-most strategy to break down intricate tasks into multiple sub-tasks, thereby improving management, efficiency and also the interpretability. An example illustrating how the Planning Agent employs this strategy to decompose a task is provided at Table 2, with the specific prompt available in Appendix C.
Task I’m planning a trip to Turkey and need information about postal codes in Istanbul. Can you provide me with the postal code and district for Istanbul province with plate number 34? Additionally, I would like to know if there are any transit agencies available in Istanbul. Please fetch their names and contact numbers.
Decomposed Sub-tasks 1. Determine the postal code and district for Istanbul province with plate number 34.
2. Find out if there are any transit agencies in Istanbul.
3. Get the names of the transit agencies in Istanbul.
4. Obtain the contact numbers for the transit agencies in Istanbul.
Table 2: Example of task decomposition by Planning Agent.

### v3

@@ Abstract
The emergence of large language models (LLMs) has opened up unprecedented possibilities for automating complex tasks that are often comparable to human performance. Despite their capabilities, LLMs still encounter difficulties in completing tasks that require high levels of accuracy and complexity due to their inherent limitations in handling multifaceted problems single-handedly. This paper introduces ‘Smurfs’, a cutting-edge multi-agent framework designed to revolutionize the application of LLMs. By seamlessly transforming a conventional LLM into a synergistic multi-agent ensemble, Smurfs can enhance the model’s ability to solve complex tasks at no additional cost. This is achieved through innovative prompting strategies that allocate distinct roles within the model, thereby facilitating collaboration among specialized agents and forming an intelligent multi-agent system. Our empirical investigation on both open-ended task of StableToolBench and closed-ended task on HotpotQA showcases Smurfs’ superior capability in intricate tool utilization scenarios. Notably, Smurfs outmatches all the baseline methods in both experiments, setting new state-of-the-art performance. Furthermore, through comprehensive ablation studies, we dissect the contribution of the core components of the multi-agent framework to its overall efficacy. This not only verifies the effectiveness of the framework, but also sets a route for future exploration of multi-agent LLM systems. Our code is available at https://github.com/FreedomIntelligence/Smurfs/.
@@ 1 Introduction
Tool manipulation has traditionally been seen as a distinctive human characteristic, dating back approximately 2.5 million years Oakley and Museum (1972); Ambrose (2001). For large language models (LLMs), access to external tools can equip them with broader capabilities beyond their fixed language modeling knowledge. For example, the search engine API empowers ChatGPT to access real-time information Zhao et al. (2023). However, LLMs still encounter several challenges when using multiple tools to solve tasks. These challenges include effective solution planning and adaptability to new tools. Hao et al. (2024); Guu et al. (2020); Qin et al. (2024).
@@
This paper addresses the critical research problem of enhancing the problem-solving capabilities of LLMs through the adoption of a plug-and-play multi-agent system (MAS) framework Dorri et al. (2018); Van der Hoek and Wooldridge (2008). We posit that a MAS approach can significantly augment the efficacy of LLMs in handling tasks that require a high degree of precision, adaptability, and comprehensive knowledge integration.
Pass Rate $\uparrow$ (%) Win Rate $\uparrow$ (%) # of Tokens per request $\downarrow$ # of Tokens per query $\downarrow$
ReACT $44.4_{\pm 1.1}$ base 1,424 6,479
DFSDT $55.4_{\pm 2.0}$ 60.4 1,743 20,714
Smurfs (ours) $57.4_{\pm 1.1}$ 62.4 459 8,096
Table 1: Comparison of token cost and performance between tool planning methods over StableToolBench. Existing methods, ReACT and DFSDT, have limitations due to high token costs or poor performance. The results are averaged over the subtasks within StableToolBench.
@@
To this end, we introduce ‘Smurfs’ an innovative MAS framework inspired by the collaborative and versatile nature of its namesake cartoon characters. The proposed framework is based on the principle: synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs. Each agent within the Smurfs framework is designed to perform specific sub-tasks, facilitating a more nuanced and effective approach to complex problem-solving. Our research delves into the architectural design, coordination mechanisms, and the operational dynamics of integrating specialized agents into a cohesive system. The effectiveness of Smurfs is validated through both open-ended and closed-ended tool planning benchmark experiments Guo et al. (2024); Yang et al. (2018), where the proposed MAS system consistently outperform baseline methods on both benchmarks. An ablation study followed by a case study further investigates the underlying reasons for this effectiveness. These results not only establish a new state-of-the-art in the field but also offer concrete evidence of the multi-agent approach’s efficacy in enhancing LLM capabilities.
The contributions of this paper can be summarized as follows:
  1. 1.
    We introduce a novel plug-and-play MAS framework to enhance the tool planning capabilities of LLMs. Experiments demonstrate the effectiveness of this approach, which is also more cost-efficient compared to existing tool planning methods.
  2. 2.
    Ablation studies further reveal the underlying reasons for the effectiveness of the MAS framework, providing valuable insights for future research.
@@ 2 Motivation
2.1 Multi-Tool Planning
To augment LLMs to do multi-tool planning for solving complex problems, previous work has seen numerous attempts. Chain-of-Thought  Wei et al. (2023) was the first to propose the method of thought and answer chain reasoning. ReACT Yao et al. (2022) further introduced the thought-action-observation format for tool chain reasoning, leading to the development of various multi-tool planning methods Chen et al. (2023a); Xu et al. (2023); Shinn et al. (2023). The latest work, DFSDT Qin et al. (2024), was proposed to address the inherent limitations of CoT and ReACT: error propagation and limited exploration. Deep First Search Decision Tree, denoted as DFSDT, is powerful in addressing multi-tool planning problems. Its core concept involves employing a depth-first search (DFS) approach for multi-tool planning (for more details, see Appendix A). When a tool fails or is deemed inadequate for solving the current problem, DFSDT backtracks to the previous solution state and attempts to resolve the issue using a different tool. However, several limitations were identified with the mechanism of DFSDT: (1) instability of the rollback mechanism, (2) redundant context, and (3) premature termination. The following sections will introduce these limitations in detail.
2.1.1 Instability of the Rollback Mechanism
The rollback mechanism in DFSDT is determined by the model. The number of steps to roll back and the selection of new tools after rollback are guided using prompt containing the errors encountered in the previous failed trajectory. When the model is sufficiently robust, this rollback mechanism serves as a highly flexible and efficient planning strategy. However, when the model’s capability is insufficient, it will fail to execute the correct rollback mechanism, i.e. retry the same error tools or roll back too far.
2.1.2 Redundant Context
In the process of planning with DFSDT, each tool plan is generated using the entire conversation history (including all the thoughts, actions, action inputs and tool responses) as context. However, in reality, each step of tool planning only requires a very small portion of the relevant history for effective planning.
@@ 2.1.2 Redundant Context
The context redundancy not only increases computational overhead but also reduces the accuracy of model inference due to the inclusion of irrelevant historical data. As highlighted by Liu et al. (2024), redundant context become particularly noticeable in tasks requiring assimilation and processing of large inputs, like verbose tool documents and API responses. The situation worsens when LLMs are supplemented with external information, such as document retrieval or online searching Petroni et al. (2020); Ram et al. (2023); Mallen et al. (2022). Although numerous language models capable of handling larger contexts are emerging Dai et al. (2019); Dao et al. (2022), they often face significant performance degradation when the important information is located at some positions Liu et al. (2024); Shi et al. (2023), which is known as the ‘lost-in-the-middle’ problem.
2.1.3 Premature Termination
The termination mechanism set by DFSDT involves adding a termination tool to the model’s selectable toolkit. When the model selects this termination tool, DFSDT stops and provides an answer. However, in practical applications, this mechanism often prematurely terminates when dealing with complex problems requiring multi-step reasoning. We hypothesize that this issue arises due to the redundant interference of other tool information and history information, which disrupts the model’s ability to judge whether the original problem should be terminated. Instead, the model focuses on whether the current sub-problem requires termination, leading the mechanism to terminate after resolving the sub-problem.
2.2 Multi Agent System
Method Multi-Agent Training Generality Reflection Planning
  ReAct Yao et al. (2022) Iterative
Reflexion Shinn et al. (2023) Iterative
Chameleon Lu et al. (2023) Global
HuggingGPT Shen et al. (2023) Global
BOLAA Liu et al. (2023) Iterative
AgentVerse Chen et al. (2023b) Iterative
FireAct Chen et al. (2023a) Iterative
DFSDT Qin et al. (2024) Iterative
RestGPT Song et al. (2023) Iterative
Lumos Yin et al. (2024) Iterative or Global
AutoAct Qiao et al. (2024) Iterative
Smurfs (Ours) Iterative and Global
Table 2: Comparison of related works.
To address the limitations inherent in DFSDT and to further enhance LLM’s multi-tool planning capabilities, multi-agent system (MAS) has emerged as a natural solution. Inspired by human social division of labor and cooperation, MAS aim to enable AI agents to accomplish more complex tasks more effectively and efficiently through the division of labor and collaboration. Previous works Song et al. (2023); Liu et al. (2023); Chen et al. (2023b); Yin et al. (2024); Qiao et al. (2024) has leveraged MAS to achieve this goal. Table 2 shows the difference between them. Based on those works, we further design the MAS named Smurfs to address issues with DFSDT. By dividing tasks among different agents, each agent can focus on a specific part of the DFSDT task, accessing only the necessary history as context during task execution, which effectively addresses the issue of redundant context. The redesign of the rollback mechanism to incorporate memory and tool list rollback mechanisms addresses the instability of the rollback mechanism. Drawing on the concept of least-to-most prompting Zhou et al. (2023), the original problem is first decomposed into sub-problems for macro-level planning. Subsequently, DFSDT is used to solve each sub-problem at the micro-level, with macro-level planning guiding the micro-level planning, thereby resolving the issue of premature termination.
@@ 3.1 Framework Overview
Figure 1 illustrates the entire workflow for the Smurfs framework. Initially, the Planning Agent identifies the user’s complex request and breaks it down into manageable sub-tasks. Executor Agents are then tasked with collecting task specific information, utilizing access to external tools. Answer Agent compiles the findings into a cohesive response, which is subsequently verified by the Verifier Agent to ensure accuracy and relevance. Each agent focus on its own task and only use the relevant part of the conversation history to reduce the Redundant Context. This process exemplifies the framework’s capability to efficiently handle complex queries by leveraging the specialized roles of multiple agents, thereby ensuring both the precision of task execution and the quality of the output. In the following sections, the system mechanism and functions of each agent will be detailed. More details of memory system can be seen at Appendix B.
3.2 Agent Components
Tools
The tool documents about the tools that Smurfs can utilize in the completion of a complex task are denoted as $D=\{n_{i},d_{i},p_{i}\}^{|d|}_{i=1}$, where n represents the tool name, d represents tool usage description, p represents parameter description and $|d|$ represents the amount of the available tools. The available tool list is denoted as $\tau=\{n_{i},d_{i}\}^{|\tau|}_{i=1}$. $\tau_{t}$ denotes the tool list Smurfs can utilize at time t.
Memory
The memory of the agent system at time t is the history of the task-solving process before t, denoted as $M=(m_{1},m_{2},...,m_{t-1})$ and $m_{i}=(\gamma_{i},a_{i})$, where $m_{i}$ represents memory element at time i and $\gamma_{i}$, $a_{i}$ represents thought and answer generated by the system at time i. There are two types of memory in Smurfs: local memory and global memory. the local memory is used to record the ongoing solution trajectory and to generate the next action in the current trajectory. The global memory, meanwhile, records all trajectories and is used to generate the sub-problem’s answer by combining all trajectory records when the maximum number of retries is exceeded. This local-global combined memory system ensures that the planning of the current solution trajectory is not influenced by the context of erroneous trajectories. It also generates an answer that combines all trajectories when the verifier agent cannot determine task completion within the maximum number of planned steps. This memory system ensures context efficiency during the task-solving process.
3.3 Macro Planning
@@
The primary responsibility of the Planning Agent is doing macro-level planning through task decomposition to prevent premature termination. The inference process of the Planning Agent is:
$Plan\>P:(p1,p2,...)=PA(q)$ (1)
Where $p_{i}$ represents sub-problem of the original query q, PA represents the Planning Agent. After the task decomposition, the agent system will use Executor Agent, Answer Agent an Verifier Agent to solve each sub-problem using DFSDT collaboratively in a sequential order. To utilize the answer of the previous sub-problem when solving subsequent sub-problem, the strategy known as least-to-most prompting Zhou et al. (2023) is used.
3.4 Subtask Solving Process
After introducing the function of plan agent, this section outlines how the agents collaborate to solve sub-tasks, as shown in Figure 2.
Stable Rollback
To address the instability of the rollback mechanism in DFSDT, we propose a rollback mechanism based on rules. Whenever an error occurs while using a tool $\tau_{t,i}$ at time t, the tool list at t $\tau_{t}$ will pop $\tau_{t,i}$ out and reperform tool selection and tool planning (ensuring that the faulty tool is not selected again). If, at time t, the tool list becomes empty, it signifies that after the system choosing tool $\tau_{t-1,j}$ at time t-1, no subsequent trajectory can solve the problem. In this case, the agent system will roll back to time t-1, meaning that the local memory M will pop out the memory element $m_{t-1}$ at time t-1, and the tool list at time t-1 $\tau_{t-1}$ will pop out tool $\tau_{t-1,j}$. The agent system will then set the time t=t-1 and continue planning. This rule-based rollback mechanism, compared to the original model-based rollback mechanism of DFSDT, is less flexible and might reduce rollback efficiency. However, it is more stable, ensuring the correctness of deep first search and enabling models with weaker capabilities to utilize DFSDT on tool planning.
Backbone Method StableToolBench
I1-Inst. I1-Cat. I1-Tool. I2-Cat. I2-Inst. I3-Inst. Average
Pass Win Pass Win Pass Win Pass Win Pass Win Pass Win Pass Win
GPT-3.5 Turbo ReACT $41.6_{\pm 1.2}$ / $48.4_{\pm 0.5}$ / $52.5_{\pm 0.5}$ / $52.2_{\pm 1.0}$ / $31.6_{\pm 1.2}$ / $39.9_{\pm 2.0}$ / $44.4_{\pm 1.1}$ ///
GPT-3.5 Turbo DFSDT $54.1_{\pm 1.0}$ $64.4$ $60.1_{\pm 0.0}$ $61.4$ $59.9_{\pm 1.7}$ $53.8$ $60.9_{\pm 0.9}$ $62.9$ $52.8_{\pm 3.7}$ $66.0$ $44.3_{\pm 4.8}$ $54.1$ $55.4_{\pm 2.0}$ $60.4$
GPT-3.5 Turbo Smurfs $\underline{60.3_{\pm 1.5}}$ $65.0$ $67.0_{\pm 1.0}$ $\underline{69.9}$ $60.3_{\pm 1.3}$ $54.4$ $54.3_{\pm 0.4}$ $\underline{63.7}$ $42.6_{\pm 1.6}$ $64.2$ $60.1_{\pm 1.0}$ $57.4$ $57.4_{\pm 1.1}$ $62.4$
Mistral-7B ReACT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Mistral-7B DFSDT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Mistral-7B Smurfs $\mathbf{76.3_{\pm 0.8}}$ 63.8 $\mathbf{86.7_{\pm 1.2}}$ 62.7 $\mathbf{81.0_{\pm 1.9}}$ 58.2 $\mathbf{70.4_{\pm 2.7}}$ 54.0 $\mathbf{63.8_{\pm 2.4}}$ $\underline{67.0}$ $\mathbf{85.2_{\pm 0.7}}$ $57.4$ $\mathbf{77.2_{\pm 1.6}}$ $60.5$
GPT-4 Turbo ReACT $41.1_{\pm 1.5}$ $60.1$ $53.2_{\pm 1.3}$ $62.1$ $42.2_{\pm 1.1}$ $48.1$ $50.0_{\pm 0.7}$ $57.3$ $38.7_{\pm 0.8}$ $65.1$ $37.7_{\pm 1.3}$ $47.5$ $43.8_{\pm 1.1}$ $56.7$
GPT-4 Turbo DFSDT $52.7_{\pm 1.4}$ $\underline{69.9}$ $58.2_{\pm 0.9}$ $66.0$ $59.7_{\pm 1.2}$ $\underline{58.2}$ $59.3_{\pm 0.7}$ $62.1$ $52.2_{\pm 2.3}$ $\mathbf{67.9}$ $61.5_{\pm 1.8}$ $\underline{65.6}$ $57.3_{\pm 1.4}$ $\underline{65.0}$
GPT-4 Turbo Smurfs $59.3_{\pm 1.4}$ $\mathbf{71.2}$ $\underline{73.3_{\pm 1.3}}$ $\mathbf{72.5}$ $\underline{67.4_{\pm 0.7}}$ $\mathbf{69.6}$ $\underline{66.7_{\pm 1.9}}$ $\mathbf{73.4}$ $\underline{55.5_{\pm 1.4}}$ $66.0$ $\underline{70.5_{\pm 0.0}}$ $\mathbf{72.1}$ $\underline{65.5_{\pm 1.1}}$ $\mathbf{70.8}$
Table 3: The open-end tool planning task evaluation on the StableToolBench benchmark Guo et al. (2024). The most effective approach is highlighted in bold, while the second-best is underlined. Win rate is calculated by comparing each model with ChatGPT-ReACT. A win rate higher than 50% means the model performs better than ChatGPT-ReACT.